OpenStudy (anonymous):

6b-12=-48 ugh.?im terrible at these kinna problems please help me find out what the b is

ganeshie8 (ganeshie8):

to find the value of \(b\), you need to isolate it

ganeshie8 (ganeshie8):

start by adding \(12\) to both sides

ganeshie8 (ganeshie8):

6b - 12 = -48 +12 +12

OpenStudy (anonymous):

Ok I got that part

ganeshie8 (ganeshie8):

good, notice that on left side, -12 and +12 eat each other out and disappear

ganeshie8 (ganeshie8):

you get : 6b = -36

OpenStudy (anonymous):

Ok thank you so much!

ganeshie8 (ganeshie8):

we're not done yet

ganeshie8 (ganeshie8):

\(b\) was not isolated fully, still the \(6\) is sticking to it, get rid of it

ganeshie8 (ganeshie8):

divide \(6\) both sides

ganeshie8 (ganeshie8):

6b = -36 /6 /6

ganeshie8 (ganeshie8):

b = -6

OpenStudy (anonymous):

Well now I know is to work the problem so thanks again :)

ganeshie8 (ganeshie8):

good to hear :) you're welcome !